### v4

# CAOTE: KV Cache Eviction for LLMs via Attention Output Error-Based Token Selection

Abstract
While long context support of large language models has extended their abilities, it also incurs challenges in memory and compute which becomes crucial bottlenecks in resource-restricted devices. Token eviction, a widely adopted post-training methodology designed to alleviate the bottlenecks by evicting less important tokens from the cache, typically uses attention scores as proxy metrics for token importance. However, one major limitation of attention score as a token-wise importance metrics is that it lacks the information about contribution of tokens to the attention output. In this paper, we propose a simple eviction criterion based on the contribution of cached tokens to attention outputs. Our method, CAOTE, optimizes for error due to token eviction, by seamlessly integrating attention scores and value vectors. This is the first method to use information from the value vector on top of attention-based eviction scores. Additionally, CAOTE can act as a meta-heuristic method with flexible usage with any token eviction method. We show that CAOTE, when combined with state-of-the-art attention score-based methods, always improves accuracies on the downstream task for LLAMA3 and QWEN2.5 model families, indicating the importance of leveraging information from values during token eviction process.
Machine Learning, ICML
token eviction, training-free long context
1 Introduction
Large Language Models (LLMs) represent a large step forward in natural language processing by demonstrating remarkable proficiency in tasks such as text generation (Coenen et al., 2021), machine translation (Xiao et al., 2023), and question-answering (Robinson et al., 2022). Many of these tasks require handling long prompt inputs efficiently, particularly in retrieval-augmented generation (RAG), long-form document understanding (Liao et al., 2024), summarization (Zhang et al., 2024a), and multi-turn dialogue systems (Thoppilan et al., 2022)—collectively referred to as long-context LLMs. A key challenge in long-context applications is the increased latency during inference and generation, primarily due to the quadratic computational complexity of self-attention and the growing memory demands of handling long sequences.
To alleviate computational overhead, Key-Value (KV) caching is a widely adopted technique that enables faster inference in LLMs. It does so by storing the key-value states of previously processed tokens and reusing them when generating new tokens, reducing redundant computations (Pope et al., 2023). However, while KV caching significantly enhances efficiency, it comes at the cost of substantial memory consumption as mentioned in (Zhang et al., 2024b) , especially in long-context setups. In such cases, the memory footprint of the KV cache often exceeds the memory usage of the model itself, making it a major bottleneck, especially, for memory-constraint devices. In this paper we focus on improving KV cache eviction methods by keeping relevant KV to utilize the memory budget in the best possible way.
Several methods have been proposed to optimize KV cache memory usage, including sparse attention mechanisms that selectively attend to a subset of tokens rather than all previous tokens reducing memory and computational complexity (Zhang et al., 2023; Xiao et al., 2024), efficient attention architectures such as linearized attention (Choromanski et al., 2020), and memory-efficient transformers (Shazeer, 2019) which approximate self-attention to minimize memory consumption. These methods require from-scratch training or model finetuning, however, in this work, we focus on post-training token eviction method which dynamically discard less important tokens from the KV cache to control memory growth (Zhang et al., 2024b). Token eviction methods offer a distinct advantage in memory management by explicitly reducing the KV cache size while preserving critical information. These methods leverage the sparsity of attention — a phenomenon where a small subset of cached keys contribute disproportionately in the attention mechanism — to selectively retain only the most important tokens.
Typically, token importance is determined based on attention scores or, equivalently, attention weights, which measure the alignment between query and key tokens (Zhang et al., 2024b; Oren et al., 2024). However, since the attention output is a linear combination of attention weights and value vectors, evicting tokens simply based on attention scores may lead to suboptimal decisions, as it does not consider the contribution of value vectors from the tokens to be evicted.
In this paper, we propose CAOTE, a post-training cache eviction method that seamlessly integrates eviction (attention) scores with value vectors by optimizing the eviction error, unlike recent methods, as shown in Table 1. CAOTE offers flexibility by being applicable to any post-training eviction method and also includes an efficient counterpart. When combined with recent cache eviction methods (Zhang et al., 2024b; Oren et al., 2024; Li et al., 2024), we observe performance boost for all recent token eviction methods on a variety of downstream tasks: 16 tasks from LongBench, Needle-in-Haystack and Perplexity measurement tasks.
The paper is divided into the following sections: Section 2 discusses background for token eviction. Our main method is in Section 3 which consists of CAOTE. In Section 4, we present experiments and results. Section 5 consists related works and Section 6 consists the conclusion.
Table 1: Overview of recent token eviction methods compared to CAOTE based on components used during eviction.
Method Keys Values Minimize eviction error
H2O
TOVA
SnapKV
X+CAOTE
2 Background
Token eviction is a popular methodology (Xiao et al., 2024; Oren et al., 2024; Zhang et al., 2024b) for decoder-only transformer inferences that prevents KV-Cache from growing linearly as token generation continues by preventing less important tokens from being cached. This has dual benefits; first, it limits memory consumption for the KV-cache and second, it reduces the computational complexity of the attention mechanism. Here, we consider the case of processing the input prompt block-wise in resource-restricted environments. In this case, token eviction can save memory and computation not only in the generation phase, but also in the prefill phase, leading to a shorter time-to-first-token which is especially beneficial when the input prompt is extremely long.
With a sequence of hidden states $X^{l}=[x_{1}^{l},\dots x_{t}^{l}]\in\mathbb{R}^{t\times d}$, the transformer block updated the hidden states as follows:
$X^{l+1}=\Phi_{\text{TRANS}}^{l}(X^{l})=\phi^{l}_{\text{FF}}\left(\phi^{l}_{\text{SA}}(X^{l})\right)$ (1)
where, $x^{l}_{j}$ is the hidden state of token $j$, $\phi_{FF}^{l}$ denotes the feedforward layer, and $\phi_{SA}^{l}$ denotes the self-attention layer, superscript $l$ denotes the layer index.
For brevity, we omit normalization layers and skip connections.
Prompt prefill
Given hidden-states $X^{l}\in\mathbb{R}^{t\times d}$ of $t$ tokens, the self-attention layer process inputs as follows:
$X^{l}_{\text{attn}}=\phi_{\text{sa}}(Q^{l},K^{l},V^{l})=\underbrace{\text{Softmax}(Q^{l}(K^{l})^{\top})}_{A^{l}}V^{l}$ (2)
where, $Q^{l},K^{l},V^{l}\in\mathbb{R}^{t\times d}$ and $A^{l}\in\mathbb{R}^{t\times t}$. Here, we omit output layer projection and multi-head extention for brevity.
Block-wise prompt prefill
Instead of processing all tokens at once (resulting in attention matrix: $A^{l}\in\mathbb{R}^{t\times t}$), we can process tokens in block-size $m$, which also helps in evicting tokens in small blocks instead of larger chunks.
$X^{l}_{\text{attn},t+1:t+m}=$ (3)
$\underbrace{\text{Softmax}(Q^{l}_{t+1:t+m}[K^{l}_{:t},\mathbf{K^{l}_{t+1:t+m}}]^{T})}_{A^{l}\in\mathbb{R}^{m\times(t+m)}}[V^{l}_{:t},\mathbf{V^{l}_{t+1,(t+m)}}]$
where, the new token hidden states are $X_{t+1:t+m}^{l}$ which are projected to $Q^{l}_{t+1:t+m},K^{l}_{t+1:t+m},V^{l}_{t+1:t+m}$
Generation. In autoregressive generation a single token is generated at each iteration
$X^{l}_{\text{attn},t+1}=\underbrace{\text{Softmax}(Q^{l}_{t+1}[K^{l}_{:t},\mathbf{K^{l}_{t+1}}]^{T})}_{A^{l}\in\mathbb{R}^{1\times(t+1)}}[V^{l}_{:t},\mathbf{V^{l}_{t+1}}]$ (4)
For resource-constraint hardware, single-inference KV cache prefill for a large number of input tokens may cause out-of-memory error or slow throughput. On the other hand, combining block-wise prefill with token eviction after processing each block of prompt can resolve this issue and improve throughput (Holmes et al., 2024; Agrawal et al., 2023). For a block-size $m$, when $b$ tokens are initially processed, the usage of memory and computation power can always be kept within budget constraints by processing the next $m$ tokens and evicting the next $m$ tokens. In this case, attention matrix has size: $A^{l}\in\mathbb{R}^{m\times(b+m)}$.
Recent eviction methods use variants of attention scores from $A^{l}$ for evicting tokens by using a function (or operator) to map $f_{\text{score}}(A^{l}):\mathbb{R}^{m\times(b+m)}\to\mathbb{R}^{b+m}$, where $f_{\text{score},j}$ is the retention score (or score) for token $j$, the top-$b$ tokens are retained based on the score: $\underset{b}{\operatorname*{argmax}}\;f_{\text{score}}(A^{l})$, where $b$ is the budget (maximum tokens allowed per layer). Examples of score functions, for H2O, $f_{\text{score},j}=\Sigma_{i=1}^{m}A_{i,j}$, and for TOVA, $f_{\text{score},j}=A^{l}_{-1,j}$. The process of token eviction follows intuitive steps as shown: computing scores for newly processed tokens, choosing top-$b$ tokens, computing attention output using the top-$b$ tokens’ hidden-state, we show the steps below:
$A^{l}_{b+m}=$ $\text{Softmax}(Q^{l}_{b+1:b+m}[K^{l}_{:b},\mathbf{K^{l}_{b+1:b+m}}]^{T})$ (5)
$i_{1},\dots,i_{b}=$ $\underset{j\in\{1,\dots,b\}}{\operatorname*{argmax}}f_{\text{score},j}(A^{l}_{b+m})$ (6)
$X^{l}_{\text{attn}}=$ $\text{Softmax}(Q^{l}_{b+1:b+m}(K^{l}_{i_{1}:i_{b}})^{T})V^{l}_{i_{1}:i_{b}}$ (7)
where the key in bold are correspond to the new tokens’ hidden states being inserted. In above equation we assume that no new query token was evicted for ease of notation. During generation, the flow remains same with $m=1$.
3 CAOTE: KV Caching through Attention Output-Based Token Eviction
Our method is developed based on two key insights: (i) existing token eviction policies primarily rely on attention scores derived from queries and keys, and (ii) attention output is a linear combination of values. We find that optimizing for eviction error is same as change in attention output due to eviction, which can be computed in closed-form for each token during generation and can be used as the eviction score (CAOTE score).
Figure 1: General flow of cache eviction when CAOTE is integrated with existing cache eviction methods. We compute the impact of removal of each token to the attention output, this is same as eviction error (or CAOTE score: $c_{1},c_{2},\dots c_{n+1}$). The token with the least impact is removed.
We first introduce CAOTE in Subsection 3.1 and how to compute eviction error in closed-form. This is followed by a discussion of its meta-property, demonstrating its applicability with other score-based attention methods such as H2O (Zhang et al., 2024b) in Subsection 3.2. Finally, we propose an efficient approximation of CAOTE in Subsection 3.3. The general workflow of CAOTE is illustrated in Fig. 1, highlighting that the modifications to existing token eviction methods are minimal.
3.1 CAOTE Score
The objective of our token eviction is to minimize eviction error: the change in attention output before and after eviction. We formulate eviction error for the generation scenario in which a single new token is inputted and therefore a single token needs to be evicted to maintain the budget $b$. Throughout the paper, we will use eviction error and CAOTE score interchangeably.
Given the attention scores of $b+1$ tokens $A=[\alpha_{1},\dots\alpha_{b+1}]\in\mathbb{R}^{1\times b+1}$ w.r.t. the last input token and the values: $V=[v_{1},\dots,v_{b+1}]\in\mathbb{R}^{d_{\text{head}}\times b+1}$, where $d_{\text{head}}$ is the head dimension. The CAOTE score for token $j\in\{1,\dots,b+1\}$ is defined as (we ignore the layer and head dependence for simplicity).
[…]
We proof that CAOTE score is same as the eviction error. We define eviction error for token $j$ as the mean square error between attention output before and after eviction. Using the same setup as above:
[…]
where, $X_{\text{attn}}$ is attention output before eviction and $X_{\text{attn},j}^{\prime}$ is attention output after eviction token $j$.
$X_{\text{attn}}=$ $\,\alpha_{1}v_{1}+\alpha_{2}v_{2}+\dots\alpha_{b+1}v_{b+1}=VA^{T}$ (10)
$X_{\text{attn},j}^{\prime}=$ $\,\alpha_{1}^{\prime}v_{1}+\dots\alpha_{j-1}^{\prime}v_{j-1}$ (11)
$+\alpha_{j+1}^{\prime}v_{j+1}\dots\alpha_{b+1}^{\prime}v_{b+1}$
where, $\alpha_{i}^{\prime}\,\forall i\in\{1,\dots,j-1,j+1,\dots b+1\}$ in Eq. (11) is the post-eviction attention score to maintain the sum of the attention score property of sum equal to $1$. In the following we show the relation between the pre and post eviction attention score for token $i$ after the eviction of token $j$.
Theorem 3.1.
Given a new input token that exceeds the budget ($b$) by $1$. A token needs to be evicted. For any token $j$ being evicted, given the retention scores pre-eviction and post-eviction for any token $i\neq j$ as $\alpha_{i}$ and $\alpha_{i}^{\prime}$ respectively, then the following relation holds:
$\alpha_{i}^{\prime}=\frac{\alpha_{i}}{1-\alpha_{j}}$ (12)
Proof.
Let the last input token has index $n$, then we define
$S\triangleq$ $\sum_{l=1}^{n}\exp(q_{n}^{T}k_{l})$ (13)
$S^{\prime}_{j}\triangleq$ $S-\exp(q_{n}^{T}k_{j})$ (14)
The retention score for token $i$ after evicting token $j$ is
$\alpha_{i}^{\prime}=$ $\frac{\exp(q_{n}^{T}k_{i})}{S_{j}^{\prime}}=\frac{\exp(q_{n}^{T}k_{i})}{S-\exp(q_{n}^{T}k_{j})}=\frac{\alpha_{i}}{1-\alpha_{j}}$ (15)
Theorem 3.2.
Given a new input token that exceeds the budget ($b$) by $1$. A token needs to be evicted. For any token $j$ being evicted, the eviction error from Eq. (9) and CAOTE score from Eq. (8) are exactly same:
$c_{j}=e_{\text{eviction},j}$ (16)
Proof.
Using Theorem 3.1, we can rewrite post-eviction attention output from Eq. (11)
$X_{\text{attn},j}^{\prime}=$ $\frac{1}{1-\alpha_{j}}\big{(}\alpha_{1}v_{1}+\dots+\alpha_{j-1}v_{j-1}$ (17)
$+\alpha_{j+1}v_{j+1}+\dots\alpha_{b+1}v_{b+1}\big{)}$
$=$ $\frac{1}{1-\alpha_{j}}\big{(}X_{\text{attn}}-\alpha_{j}v_{j}\big{)}$ (18)
Replacing Eq. (18) in Eq. (9), we get
[…]
$=$ $c_{j}$ (19)
Hence proved. ∎
Using Eq. (19) CAOTE scores (or eviction error) for each token can be computed in parallel as the dependency on only on attention scores and value vectors. Note that this is the firsts formulation seamlessly integrating attention scores and value vectors into a single score. Any norm can be used for computing CAOTE score and based on empirical results we choose $L_{2}$-norm. Evicting multiple tokens using CAOTE formulation is discussed in Appendix B.
3.2 CAOTE with general score-based eviction methods
The CAOTE formulation allows the use of arbitrary scoring-based eviction methods to incorporate the values into their scoring mechanism, provided that the scores sum to 1.0. In practice, we can adjust the raw eviction scores without changing their relative order by simple normalizations (affine transformations). Let $H$ be the set of retention scores and $f^{\text{norm}}$ be the normalizing function. The CAOTE score for general eviction methods is given by:
$c_{j}$ $=f^{\text{caote}}_{j}(f^{\text{norm}}(H),V)$ (20)
[…]
where, $h_{j}^{\text{norm}}=f^{\text{norm}}_{j}(H)$. We further discuss the generalization of CAOTE to well-known token eviction methods in the following.
CAOTE for H2O
We consider H2O (Zhang et al., 2024b), where the scores ($H=[h_{1},\dots,h_{b+1}]$) are based on the sum of previous attention scores, leading to $\Sigma_{j=1}^{b+1}h_{j}>1$ during generation-phase as proved in Theorem 36 in Section D.1. In this case, simply dividing each token score by the sum of all scores maps the scores to the range $[0,1]$ and ensures that new scores follow $\sum_{i=1}^{b+1}h_{i}^{\text{norm}}=1$.
$h_{j}^{\text{norm}}=\frac{h_{j}}{\Sigma_{i=1}^{b+1}h_{i}}$ (22)
For recent methods where all the scores are $\geq 0$, simply dividing by sum of all scores suffices. Note that for TOVA (Oren et al., 2024), this summation is by default equal to $1$.
3.3 Fast CAOTE Computation
We also propose a compute-efficient version of CAOTE, FastCAOTE, with negligible performance degradation and minimal compute overhead relative to CAOTE. Here, the pre-eviction attention output ($X_{\text{attn}}$) is replaced with mean of values while everything else remains same, that is, CAOTE score for token $j$ is:
[…]
We find empirically, that FastCAOTE is highly correlated to CAOTE by observing high ($\geq 0.8$) Spearman’s correlation between them for each layer as shown in Table 8.
We provide theoretical inference time latency for using CAOTE and FastCAOTE in Appendix C, where we observe that FastCAOTE leads to minimal increase in latency.
4 Results
In this section, we demonstrate the efficacy of CAOTE for boosting performance on state-of-the-art token eviction methods on a wide range of downstream benchmarks.
4.1 Experiment Setup
Tasks
We study the impact of CAOTE on different token eviction methods by evaluating on LongBench (Bai et al., 2024), covering single QA, multiple QA, single/multi-document summarization, synthetic, and code generation tasks. We measure long-context perplexity on the Booksum dataset (Kryściński et al., 2021), and lastly, measure recall accuracy on Needle In A Haystack task (Liu et al., 2024; Kamradt, 2023).
Baselines
We compare the performance of CAOTE to various token eviction methods including: H2O (Zhang et al., 2024b), TOVA (Oren et al., 2024), and SnapKV (Li et al., 2024), on LLAMA3 models: Llama 3.2-3B-Instruct and Llama 3.1-8B-Instruct (Dubey et al., 2024), and QWEN2.5 models: Qwen 2.5-3B-Instruct and Qwen 2.5-7B-Instruct (Yang et al., 2024) for all subsequent experiments.
Budgets
We evaluated all methods with various KV cache budget sizes of $2048$, $4096$, $6144$, and $8192$, denoted by 22k, 44k, 66k, and 88k, respectively.
Prompt consumption
Unlike other token eviction methods that assume to prefill prompt at once followed by KV cache eviction, we propose to consume tokens in block-wise manner as described in Section 2 with the block-size of $128$, i.e., at each inference of LLM during the prefill phase, there are $128$ new tokens incoming and being added to the cache, and $128$ tokens from the cache are evicted once the total number of tokens reaches the cache budget limit.
4.2 LongBench
Table 2: LongBench results for Llama 3.1-8B and Llama 3.2-3B-Instruct. Higher number is better. We highlight the best performing methods within a given budget with bold and the second best with underline.
Single Doc. QA Multi Doc. QA Summarization Fewshot Learning Synthetic Code
Narrative QA Qasper MF-en HotpotQA 2WikiMQA Musique GovReport QMSum MultiNews TREC TriviaQA SAMSum PCount PR-en Lcc RB-P Avg.
Llama 3.1-8B 30.05 47.00 56.12 57.33 47.81 32.25 34.86 25.32 27.02 73.00 91.61 43.37 8.33 99.50 61.66 51.94 49.20
2k H2O 1.74 21.15 25.33 26.11 24.15 8.78 2.17 2.70 16.78 44.00 29.36 7.62 2.25 5.88 40.15 12.14 16.89
+ CAOTE 14.32 38.34 45.97 37.77 42.51 22.06 29.57 15.11 27.02 62.00 63.60 27.34 2.00 15.50 56.99 32.87 33.31
+ FastCAOTE 15.15 41.27 46.6 39.91 40.02 24.55 30.05 16.19 26.95 63 62.39 26.86 3.08 17.5 56.87 34.75 34.07
TOVA 22.57 37.26 39.43 45.74 34.48 14.77 28.87 21.17 26.95 62.50 90.73 42.74 0.00 18.00 62.68 52.48 37.52
+ CAOTE 21.92 37.47 38.28 45.88 35.2 15 29.02 21.21 27 62.5 91.34 43.22 1.5 23 62.6 54.13 38.08
+ FastCAOTE 21.94 38.22 38.22 46.72 36.93 14.31 29.06 21.72 26.98 63 91.65 43.53 1.5 22 62.44 52.88 38.19
SnapKV 21.81 37.22 37.19 46.10 35.42 16.53 29.83 21.05 26.77 61.00 88.84 42.56 4.03 51.50 62.37 51.45 39.60
+ CAOTE 21.75 37.49 36.86 44.62 37.26 16.82 30.3 21.67 26.88 64 90.65 42.8 2.09 53 62.5 52.09 40.05
+ FastCAOTE 23.26 38.54 39.16 43.2 38.27 17.54 30.28 21.97 26.76 65.5 90.91 42.71 2.84 56 62.36 52.4 40.73
4k H2O 4.07 36.16 36.00 33.52 32.87 17.78 6.66 5.95 24.09 55.00 47.65 17.41 4.00 24.50 54.85 21.43 26.37
+ CAOTE 20.28 46.08 51.45 47.38 46.05 30.89 33.39 20.8 26.93 69 80.12 38.27 4.31 32 59.22 40.51 40.42
+ FastCAOTE 24.4 44.32 48.11 48.19 43.69 21.12 31.55 22.36 26.98 65 91.18 43.11 2 46.5 61.62 53.35 42.09
TOVA 22.68 44.55 47.87 46.76 44.54 20.56 30.95 22.13 26.96 61.50 90.56 43.27 3.00 43.50 61.62 53.40 41.49
+ CAOTE 24.68 43.88 48.07 49.64 44.91 22.57 31.25 22.25 26.98 63 91.29 43.29 2.5 46.5 61.6 53.45 42.24
+ FastCAOTE 24.4 44.32 48.11 48.19 43.69 21.12 31.55 22.36 26.98 65 91.18 43.11 2 46.5 61.62 53.35 42.09
SnapKV 24.79 44.22 47.30 48.49 46.73 20.55 32.19 22.68 26.95 67.50 90.98 43.14 5.17 89.50 61.44 51.20 45.18
+ CAOTE 24.41 43.16 47.77 50.87 44.11 21.04 32.51 22.98 26.93 69 91.31 43.18 3.33 92 61.04 51.74 45.34
+ FastCAOTE 24.12 44.59 47.39 50.82 44.07 22.38 32.33 22.92 27.01 69 91.31 43.53 4.58 94.5 61.31 52.11 45.75
Llama 3.2-3B 23.76 40.23 50.09 50.69 42.29 26.84 33.09 24.30 25.21 72.50 90.11 42.58 3.00 96.50 56.22 56.52 45.87
2k H2O 1.63 19.96 20.20 18.02 19.56 2.88 0.78 1.55 15.97 41.00 21.97 9.83 0.50 0.50 39.71 13.91 14.25
+ CAOTE 6.38 34.36 40.6 32.52 31.08 12.69 27.36 15.04 24.6 59 52.83 26.78 3.7 7.56 51.09 36.33 28.87
+ FastCAOTE 7.27 34.23 39.74 32.22 30.08 12.63 27.86 15.48 25.15 60.5 53.09 26.94 2.17 8.12 51.2 35.06 28.86
TOVA 17.14 30.14 32.44 35.96 30.05 13.08 26.15 19.70 25.04 56.50 87.81 40.48 2.50 11.50 55.51 52.36 33.52
+ CAOTE 17.75 30.45 32.19 37.53 29.35 13.33 26.92 19.93 25.18 60.5 88.08 41.65 1.00 12.5 54.92 53.22 34.03
+ FastCAOTE 17.93 30.52 33.1 37.01 30.7 13.88 26.39 20.28 24.96 60.5 88.95 41.27 2.00 12.5 55.65 53.56 34.33
SnapKV 17.38 31.37 31.48 37.77 30.05 11.54 27.03 19.93 24.97 59.00 88.13 40.48 3.50 32.50 56.32 55.91 35.46
+CAOTE 19.11 33.12 31.09 38.68 32.09 12.31 27.48 20.38 25.2 64 87.7 40.78 2.5 35 57.03 56.21 36.42
+FastCAOTE 18.96 32.97 33.61 39.00 31.36 12.35 27.48 20.15 25.24 65 87.2 40.7 4.5 36.5 56.06 57.12 36.76
4k H2O 2.92 31.94 33.23 24.49 28.08 7.55 5.44 6.30 22.77 53.00 38.85 20.33 1.50 7.50 51.23 22.94 22.38
+CAOTE 12.87 40.79 47.56 40.28 39.07 16.61 30.82 19.65 25.12 65.5 69.29 34.16 2.35 17 55.32 45.12 35.09
+FastCAOTE 11.85 40.41 47.93 40.81 38.93 17.36 31.22 19.67 25.1 65 71.25 34.89 3.5 15 55.5 44.3 35.17
TOVA 20.52 39.53 42.47 44.12 38.42 18.22 29.36 21.36 24.96 63.50 88.98 41.50 3.00 23.50 55.72 56.66 38.24
+CAOTE 19.59 39.79 42.03 45.25 37.07 19.3 29.39 21.57 24.92 63 89.14 41.77 3.00 29.5 55.68 56.19 38.57
+FastCAOTE 19.77 39.23 43.13 45.28 37.04 18.82 29.25 21.94 24.96 63 88.64 41.92 3.5 29 55.68 56.41 38.6
SnapKV 19.85 39.22 39.86 46.70 37.98 16.64 29.79 21.21 25.01 65.50 89.35 40.95 2.50 62.50 55.74 56.88 40.60
+CAOTE 20.1 39.74 41.01 45.64 38.26 18.64 30.07 21.53 24.98 67.5 89.08 41.78 3.00 67 55.73 56.71 41.30
+FastCAOTE 19.68 39.24 41.03 44.52 39.09 18.62 30.15 21.72 24.97 67 88.86 41.24 3.00 71 55.67 56.64 41.40
Table 3: LongBench results for Qwen 2.5-7B/2.5-3B-Instruct. Higher number is better. We highlight the best performing methods within a given budget with bold and the second best with underline.
Single Doc. QA Multi Doc. QA Summarization Fewshot Learning Synthetic Code
Narrative QA Qasper MF-en HotpotQA 2WikiMQA Musique GovReport QMSum MultiNews TREC TriviaQA SAMSum PCount PR-en Lcc RB-P Avg.
Qwen 2.5-7B 15.75 16.94 32.38 11.89 11.88 7.95 34.33 19.91 22.67 65.5 87.05 44.75 4.22 93.08 57.74 61.84 36.74
2k H2O 2.39 7.29 12.42 8.55 11.06 2.73 3.62 6.6 15.69 42.5 28.21 10.63 0.65 0 35.1 18.77 12.89
+ CAOTE 4.55 14.3 27.58 11.33 13.55 7.76 26.65 15.62 22.93 57 49.78 27.74 1.54 11.08 51.45 32.7 23.47
+ FastCAOTE 4.8 12.79 28.72 12.94 13.25 7.53 27.06 14.46 22.84 59 48.23 26.4 2.53 11.54 52.85 32.93 23.62
TOVA 8.49 14.01 21.04 14 11.51 5.09 27.43 17.84 22.83 56.5 79.56 40.55 2.43 9.29 55.99 56.15 27.67
+ CAOTE 10.46 14.82 25.06 14.62 11.73 6.01 27.66 18.02 22.78 57.5 79.39 40.87 2.5 11.25 56.22 56.51 28.46
+ FastCAOTE 10.08 13.58 25.28 14.44 12.14 5.24 27.34 18.31 23.11 55.5 78.51 41.67 2.7 10.54 56.56 58.05 28.32
SnapKV 11.6 12.45 23.66 12.38 10.64 7.03 27.57 18.27 22.85 58 81.78 41.13 3.76 19.42 55.83 56.53 28.93
+ CAOTE 14.02 12.23 24.55 16.45 10.35 8.59 27.77 18.91 22.87 56 80.58 40.43 2.38 21.52 55.17 56.03 29.24
+ FastCAOTE 14.26 14.11 24.11 15.31 11.35 7.88 27.95 18.86 22.74 56.5 80.92 41.49 3.8 22.42 55.89 57.43 29.69
4k H2O 1.99 11.92 19.88 10.24 10.12 4.73 9.08 10.14 20.85 51.00 37.37 20.57 3.16 6.43 52.14 29.09 18.67
+ CAOTE 4.78 18.06 32.49 16.23 17.28 9.57 29.81 18.04 22.86 59.5 63.05 36.91 2.7 28.25 55.13 42.42 28.57
+ FastCAOTE 5.69 16.99 32.62 18.22 16.58 10.48 30.3 17.71 22.88 59.5 62.95 36.29 2.1 27.65 56.3 40.65 28.56
TOVA 12.83 17.03 27.01 16.8 13.37 8.05 29.21 19.05 22.73 58.5 82.67 42.71 1.67 15 56.69 56.59 29.99
+ CAOTE 12.97 14.99 27.53 17.94 12.93 9.21 29.76 19.7 22.92 58 82.03 43.14 2.15 17.25 57.32 59.37 30.98
+ FastCAOTE 14.52 16.71 26.97 18.73 13.84 9.59 29.47 19.45 22.87 59.5 82.96 42.42 2.6 20.33 57.22 58.42 30.98
SnapKV 14.35 13.45 28.28 16.33 11.74 8.12 29.71 19.18 22.82 57 83.8 43.27 2.41 39.83 58.12 58.67 31.69
+ CAOTE 15.07 14.34 28.7 16.7 12.89 10.54 30.03 19.58 22.73 59.5 83.12 42.56 3.17 55.92 57.34 58.85 33.19
+ FastCAOTE 17.12 14.69 27.6 17.52 13.69 9.96 30.24 20.02 22.88 58.5 81.13 42.31 4.06 53.33 57.51 58.77 33.08
Qwen 2.5-3B 18.08 22.49 39.72 27.86 20.45 18.93 32.8 23.74 24.89 67.5 85.05 43.88 5 40.97 51.91 47.53 35.68
2k H2O 1.8 9.18 11.62 8.54 7.31 2.77 5.93 6.99 16.89 38 21.87 7.69 1 3 37.36 22.9 12.68
+ CAOTE 6.9 22.71 28.09 15.23 18.19 4.95 29.53 17.68 24.74 52.5 45.81 26.95 1.92 6.16 45.81 36.48 23.98
+ FastCAOTE 7.03 22.37 28.88 15.34 16.95 5.19 29.13 18.06 25.03 54.5 46.35 25.35 2.23 7.22 45.7 36.59 24.12
TOVA 11.69 14.94 25.33 17.29 12.58 5.91 26.67 21.49 24.78 51.5 68.8 41.79 0.23 6 49.79 48.6 26.71
+ CAOTE 11.17 15.23 27.42 18.94 13.1 6.94 27.01 21.62 24.86 57.5 68.38 42.11 0.82 4.88 49.36 48.13 27.34
+ FastCAOTE 11.04 15.36 27.72 19.8 13.65 6.37 27.17 22.08 24.64 57 69.13 42.48 0.77 5.25 48.36 48.58 27.46
SnapKV 11.7 13.91 24.28 14.8 10.89 7.42 27.4 21.63 24.64 54.5 75.35 42.72 2.5 18.33 49.65 50.59 28.14
+CAOTE 12.69 14.88 26.16 13.93 12.21 7.07 27.48 20.99 24.75 61 75.58 42.08 4 21.29 49.94 52.38 29.15
+FastCAOTE 12.03 14.56 24.82 14.66 10.83 7.89 27.51 20.98 24.67 62.5 75.51 41.53 2 17 49.02 50.83 28.52
4k H2O 2.82 17.34 23.27 10.18 10.47 3.03 11.06 10.73 22.93 50.75 34.93 18.03 4.35 7.32 47.74 29.42 19.02
+CAOTE 7.63 24.16 35.29 20.17 17.67 12.61 31.14 19.04 25.01 62.5 64.84 34.19 4.25 18.37 49.79 41.45 29.26
+FastCAOTE 8.58 23.45 33.14 21.72 16.11 12.26 31.11 19.76 25.04 62 65.01 35.15 4.6 17.88 50.05 40.03 29.12
TOVA 12.19 18.31 32.56 20.58 13.8 7.74 28.82 22.27 24.98 59 80.66 43.05 1.11 9.56 49.93 46.74 29.46
+CAOTE 13.16 18.67 30.74 19.33 15.7 7.32 28.93 22.14 24.91 59.5 78.54 43.57 1.55 8.25 49.4 47.31 29.31
+FastCAOTE 12.2 18.55 32.29 19.3 15.13 7.23 29.12 22.44 24.97 60 78.8 43.12 1.5 10.25 49.6 47.74 29.52
SnapKV 12.98 2.21 31.77 18.33 14.41 10.83 29.14 22.38 24.89 61 84.17 42.63 3.75 25.42 50.22 48.77 30.18
+CAOTE 13.65 20.35 32.62 19.36 15.27 11.42 29.47 22.44 24.78 64 82.6 43 4.25 24.46 50.37 49.28 31.71
+FastCAOTE 13.46 19.92 32.53 20.44 13.64 8.44 29.56 22.14 24.93 64.5 82.73 43.26 2 24.58 50.14 50.08 31.40
We present the accuracy of Llama 3.1-8B-Instruct, Llama 3.2-3B-Instruct and, Qwen 2.5-3B-Instruct, Qwen 2.5-7B-Instruct using baseline eviction methods with budget of 22k, 44k, both with and without CAOTE in Table 2 and Table 3. We observe that the best average performance is given by SnapKV-FastCAOTE for the Llama3 models, while for Qwen 2.5 models SnapKV-CAOTE performs the best. H2O shows $30\%$ improvement with CAOTE, while TOVA, SnapKV also show overall improvements, making their average accuracy closer to dense accuracy. Additionally, for some QA tasks (Qasper, MF-en, Musique, 2WikiMQA), H2O-(Fast)CAOTE performs best. Additional results for the $6$ k, $8$ k budget are shown in Table 9, Table 10 for Llama3 and Qwen 2.5 respectively, in Section E.1, which follow a trend similar to the 22k, 44k budgets.
4.3 Perplexity
Table 4: Perplexity difference between different eviction methods with dense baseline. The lower is better. Negative entry in table means the method performs better than dense baseline. The PPL of Llama 3.2-3B-Instruct and Llama 3.1-8B-Instruct is 15.4911 and 9.833 respectively.
Budget H2O TOVA SnapKV
+CAOTE +FastCAOTE +CAOTE +FastCAOTE +CAOTE +FastCAOTE
Llama 3.1-8B-Instruct
2k 2.007 1.884 1.891 -0.046 -0.088 -0.085 -0.019 -0.097 -0.098
4k 1.284 1.079 1.061 -0.047 -0.060 -0.058 -0.0483 -0.080 -0.079
6k 0.843 0.716 0.703 -0.035 -0.0366 -0.085 -0.036 -0.043 -0.045
Llama 3.2-3B-Instruct
2k 3.814 3.561 3.563 0.493 0.442 0.432 0.555 0.451 0.435
4k 2.460 2.142 2.128 0.175 0.150 0.144 0.223 0.152 0.144
6k 1.369 1.219 1.187 0.065 0.057 0.057 0.076 0.056 0.044
8k 0.589 0.462 0.448 0.023 0.012 0.011 0.020 0.007 0.005
We use the Booksum dataset (Kryściński et al., 2021) to measure generation perplexity of different eviction methods for various budgets. In Table 4, we show perplexity gap between a model using a given eviction strategy and that of the model without token eviction with cache budgets of 22k, 44k and 66k. We observe that when CAOTE is applied to existing eviction methods, the perplexity either improves or surpasses the perplexity of the baseline model. TOVA-FastCAOTE, SnapKV-CAOTE, and SnapKV-FastCAOTE perform best for 66k, 44k, 22k budgets, respectively, for Llama 3.1-8B-Instruct; for Llama 3.2-3B-Instruct, TOVA-FastCAOTE performs best with 22k and 44k budgets and SnapKV-FastCAOTE beats other methods using 66k and 88k. Perplexity results for Qwen 2.5 models are shown in Table 11 in Section E.2.
4.4 Needle In A HayStack
(a) H2O
(b) TOVA
(c) SnapKV
(d) H2O-CAOTE
(e) TOVA-CAOTE
(f) SnapKV-CAOTE
(g) H2O-FastCAOTE
(h) TOVA-FastCAOTE
(i) SnapKV-FastCAOTE
Figure 2: Needle-In-A-Haystack accuracies of Llama 3.1-8B-Instruct with token eviction with 66k cache budget.
Table 5: Needle-in-haystack accuracy for Llama 3.2-3B/3.1-8B-Instruct using baseline eviction methods with(out) CAOTE. Higher is better, maximum accuracy is $1.0$.
Budget H2O TOVA SnapKV
+CAOTE +FastCAOTE +CAOTE +FastCAOTE +CAOTE +FastCAOTE
Llama 3.1-8B-Instruct
2k 0.174 0.270 0.264 0.196 0.204 0.202 0.214 0.226 0.242
4k 0.330 0.538 0.568 0.286 0.298 0.292 0.360 0.392 0.420
6k 0.544 0.698 0.676 0.370 0.402 0.396 0.490 0.550 0.580
Llama 3.2-3B-Instruct
2k 0.104 0.160 0.172 0.172 0.150 0.166 0.154 0.172 0.168
4k 0.198 0.262 0.294 0.220 0.232 0.232 0.226 0.222 0.232
6k 0.258 0.308 0.322 0.258 0.278 0.270 0.272 0.264 0.312
8k 0.324 0.414 0.404 0.338 0.364 0.344 0.342 0.336 0.366
Lastly, we run extensive experiments on Needle-In-A-Haystack benchmark (Liu et al., 2024; Kamradt, 2023) and show quantitative results in Table 5 and visualizations for 66k budget Llama 3.1-8B model in Figure 2. We observe in Table 5 that H2O-FastCAOTE performs best for all budgets with Llama 3.2-3B. When using a budget 44k with Llama 3.1-8B, CAOTE boosted H2O outperforms TOVA, SnapKV as well. H2O-CAOTE performs best for Llama 3.1-8B with budget = {22k, 66k} and H2O-FastCAOTE performs best for44k budget for Llama3.1-8b. The gains in precision are especially high for the $4$ k budget for the Llama 3.1-8B model, with an increase of up to $30-60\%$. We can see in Figure 2 that CAOTE improves the state-of-the-art eviction method and is able to predict beyond their budget constraints. Results for Qwen 2.5 models are shown in Table 12 in Section E.3.
5 Related Work
Sparse and Efficient Attention
Sparse or efficient attention based methods result in mitigating the computation load and saving memory consumption by using efficient linear attentions (Katharopoulos et al., 2020). Additionally, there are KV compression methods which don’t evict any tokens as post eviction the token is not retrievable, (Dong et al., 2024) proposes to keep important tokens based on attention score in cache while combining the evicted tokens via linear attention into single embedding. Landmark attention injects learnable special tokens between chunks of tokens and access past tokens in chunks instead of individually. Lastly, there are better architectures based with constant KV memory which outperform linear attention attentions (Mohtashami & Jaggi, 2023). However, all these methods require either from-scratch training or fine-tuning.
KV Cache Eviction
At the extreme end of efficient KV cache management, token eviction methods have been extensively studied. Leveraging the sparsity of attention in LLMs (Xiao et al., 2024; Sheng et al., 2023; Chen et al., 2021), these methods determine the importance of KV pairs using (learned) rules and retain the pairs with high scores in the cache to approximate the attention output. StreamingLLM (Xiao et al., 2024) observes an attention sink phenomenon, which states that the first few tokens tend to receive the majority of attention weights. To exploit this, it proposes SinkAttention, which prioritizes keeping the initial tokens in the cache while doing sliding window-based attention. Other methods, such as H2O (Zhang et al., 2024b), TOVA (Oren et al., 2024), SnapKV (Li et al., 2024), and RoCO (Ren & Zhu, 2024), retain tokens with high attention scores in the KV cache with various algorithmic modifications. These include preserving the first or last tokens in addition to those with high attention scores or applying smoothing techniques to the attention scores. While these token eviction methods primarily rely on attention scores to assess token importance, CAOTE introduces an orthogonal scoring metric that estimates the impact of values on approximating attention outputs. This metric can complement existing token importance scoring approaches, enhancing other eviction methods.
Furthermore, other lines of work enable layer-wise budget optimization (Feng et al., 2024) by considering the scores of all heads jointly, and selecting top-K nodes, while others (Ge et al., 2023) consider managing memory by keeping/discarding based on token characteristics with baseline token eviction (H2O). Our proposed method, CAOTE is highly flexible, and can be integrated with both the mentioned methods to achieve additional boost on performance.
6 Conclusion
We propose a post-training KV cache eviction method that can be seamlessly integrated with any existing eviction strategies. Our approach, CAOTE, introduces an optimization objective aimed at minimizing the alteration in attention output when evicting a token. This objective ensures the incorporation of both attention scores and value vectors in the eviction decision process. Our formulation allows for the parallel computation of the CAOTE score for all tokens. Additionally, we present an efficient variant, FastCAOTE. Through extensive evaluations across various downstream tasks, we demonstrate that eviction methods equipped with CAOTE consistently deliver superior performance.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.
Appendix A Limitation
CAOTE is a myopic (greedy) strategy and its scoring framework based on the assumption of evicting $1$ token per iteration. This assumption breaks during prefilling stage, however, taking into account change in attention output due to multi-token eviction is non-trivial. Fortunately, we observe that even assuming multi-token eviction independently (without considering the effect of other tokens being evicted), CAOTE is still able to give boost in performance for all tasks. Due to this reason CAOTE’s performance might further improve with smaller prompt filling block-size.
Appendix B Extending CAOTE to Multi-token Eviction
We start with simple example, without loss of generality consider evicting two tokens in the order $i=1,j=2$. From (8) single token eviction error for token $1$ is
[…]
After evicting token $1$, the updated attention output becomes:
$X_{\text{attn},1}^{{}^{\prime}}=\Sigma_{i=2}^{n+1}\alpha_{i}^{{}^{\prime}}v_{i},\,\text{where }\alpha_{i}^{{}^{\prime}}=\frac{\alpha_{i}}{1-\alpha_{1}}$ (25)
Now, evicting token $2$ from updated attention output above yields
$X_{\text{attn},[1,2]}^{{}^{\prime\prime}}=\Sigma_{i=3}^{n+1}\alpha_{i}^{{}^{\prime\prime}}v_{i},$ (26)
$\text{where }\alpha_{i}^{{}^{\prime\prime}}=\frac{\alpha_{i}^{{}^{\prime}}}{1-\alpha_{2}^{{}^{\prime}}}=\frac{\alpha_{i}}{1-\alpha_{1}-\alpha_{2}}$ (27)
This leads to the following insights:
  • The updated attention output can be expressed as
    $X_{\text{attn},[1,2]}^{{}^{\prime\prime}}=\frac{1}{1-\alpha_{2}^{{}^{\prime}}}(X_{\text{attn},1}^{{}^{\prime}}-\alpha_{2}^{{}^{\prime}}v_{2})$ (28)
  • Eviction error of removing token $2$ after removing token $1$ is
[…]
  • Substituting the expressions above, we obtain a closed-form joint eviction score:
[…]
This formulation generalizes to evicting $m$ tokens jointly (assuming first $m$ tokens here without loss of generality)
[…]
This expression highlights the combinatorial nature of multi-token eviction: for $n$ tokens with $m$-token eviction, there are $\binom{n}{m}$ possible combinations. This makes exact computation intractable for large $m$, motivating the need for approximate or greedy strategies.
Importantly, existing token eviction methods typically rely only on attention scores and assume independence between tokens. As a result, the relative ranking of tokens remains unchanged even after evictions, which limits their effectiveness in multi-token settings.
In contrast, CAOTE explicitly models the interplay between attention scores and value vectors, capturing how evicting one token affects the contribution of others. This introduces interdependencies between tokens during eviction, which are critical for accurate multi-token decisions.
Appendix C Inference Latency
We analyze the computational overhead of CAOTE and FastCAOTE during both prefill and generation phases. The overhead is minimal, especially for FastCAOTE.
Let $s$ be sequence length, $d$ the hidden dimension, $d_{KV}$ the KV hidden dimension, $d_{FFN}$ the intermediate dimension, $d_{V}$ the vocabulary size, and $L$ the number of hidden layers. The total floating-point Operation count are as follows:
$\text{prefill flops }:2Lsd(d+d_{KV}+d_{FFN}+2s+6)+dd_{V}$ (32)
$\text{generation flops }:2Ld(d+d_{KV}+d_{FFN}+2s+6)+dd_{V}$ (33)
$\text{{CAOTE} flops}:Ls(7d+3)$ (34)
$\text{{FastCAOTE} flops}:Ls(4d+3)+L$ (35)
We compute the relative overhead for Llama3.1-8B model with different sequence lengths and show the ratio between prefill (and generation) flops to CAOTE flops in Table 6, 7
Sequence Length Ratio
with CAOTE with FastCAOTE
4k 1.6e-4 8.9e-5
8k 9.25e-5 5.28e-5
32k 6.45e-5 3.69e-5
Table 6: Ratio between prefill flops and (Fast)CAOTE flops. CAOTE adds minimal overhead during prefill phase.
Sequence Length Ratio
with CAOTE with FastCAOTE
512 0.08 0.046
1024 0.15 0.087
Table 7: Ratio between generation flops and (Fast)CAOTE flops. CAOTE adds minimal overhead during prefill phase, especially FastCAOTE
Appendix D Additional Proofs
D.1 H2O scores
Theorem D.1.
Given H2O scores for $b+1$ tokens as $h_{1},\dots,h_{b+1}]$, the summation of all $h_{i}$, $\forall i\in\{1,\dots,b+1\}$ is greater than 11
$\Sigma_{i=1}^{b+1}h_{i}>1$ (36)
Proof.
Assuming that only $b+1$ tokens are present and are propagated through the model at the same time. The causal attention mask $A\in[0,1]^{b+1\times b+1}$, will have all entries on upper triangle excluding diagonal is 0. The first token will attend to itself have attention score as 11
$A_{1,1}=$ $1$ (37)
$A_{1,>1}=$ $0$ (38)
H2O score for token $1$ is defined as the sum of attention score to token $1$ by all future tokens:
$h_{1}=A_{1,1}+A_{2,1}+\dots+A_{b+1,1}=\Sigma_{i=1}^{b+1}A_{i,1}$ (39)
In general for a token $j$, the H2O score is defined as
$h_{j}=$ $A_{j,j}+A_{j+1,j}+\dots+A_{b+1,j}$ (40)
$=$ $\underbrace{A_{1,j}+\dots+A_{j-1,j}}_{=0\text{ causal mask}}+A_{j,j}\dots+A_{b+1,j}$ (41)
$=$ $\Sigma_{i=1}^{b+1}A_{i,j}$ (42)
Using Eq. (42) and summing for all H2O scores, we get
$\Sigma_{j=1}^{b+1}h_{j}=$ $\Sigma_{j=1}^{b+1}\Sigma_{i=1}^{b+1}A_{i,j}$ (43)
$=$ $\Sigma_{i=1}^{b+1}\Sigma_{j=1}^{b+1}A_{i,j}$ (44)
$=$ $\Sigma_{i=1}^{b+1}(\underbrace{\Sigma_{j=1}^{i}A_{i,j}}_{=1\text{ due to softmax}}+\underbrace{\Sigma_{j=j+1}^{b+1}A_{i,j}}_{=0\text{ for causal mask}})$ (45)
$=$ $b+1>1$ (46)
Hence proved.
D.2 Relation of CAOTE score (eviction error) to logits
We show that evicting token based on CAOTE score can lead to smaller discrepancy in final logits which affect the downstream performance. As CAOTE score is the eviction error during generation phase, we instead show the relation between eviction error and logits. We start by showing for a single attention layer (single head) based network, its extension to multiple heads and, finally a transformer layer (self-attention and feed-forward-network). For simplicity we ignore layer-norms. Some definitions which will used are given below. We assume a budget of $b$, with current token sequence having $b+1$ tokens (superscript denotes layer):
$X^{0}\triangleq[x_{1},x_{2},\dots,x_{b+1}],\,X^{0}_{b+1}\triangleq x_{b+1}$ (47)
The attention output for a sequence of $b+1$ tokens is (for layer $m$)
$X_{A,b+1}^{m}\triangleq\Sigma_{j=1}^{b+1}\alpha_{j}^{m}W_{v}^{m}X_{j}^{m}$ (48)
The logits with and without eviction for token $j$ are defined as $l_{j},\hat{l}_{j}$ respectively.
(Case 1) Single self-attention layer with single head: The logits for dense baseline is:
$X^{1}_{b+1}=$ $X^{0}_{b+1}+W_{O}X_{A,b+1}^{0}$ (49)
$l_{b+1}=$ $W_{H}X^{1}_{b+1}=W_{H}(X^{0}_{b+1}+W_{O}\Sigma_{j}\alpha_{j}W_{V}X^{0}_{j})$ (50)
where $W_{H},W_{O},W_{V}$ are the LM-head, output projection, and value projection respectively. $X^{1}_{b+1}$ is the output after the residual connection.
The logits with eviction will have a perturbation due to error in attention output (CAOTE score or eviction error), and is given as:
$\hat{X}^{1}_{b+1}=$ $X^{0}_{b+1}+W_{O}X_{A,b+1}^{0}+{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}W_{O}\Delta_{A}^{0}}$ (51)
$\hat{l}_{b+1}=$ $W_{H}\hat{X}_{b+1}^{1}$ (52)
$=$ $W_{H}(X^{0}_{b+1}+W_{O}X_{A,b+1}+{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}W_{O}\Delta_{A}^{0}})$ (53)
$=$ $l_{b+1}+{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\Delta_{l,b+1}}$ (54)
where the logit error is $\Delta_{l,b+1}=W_{H}W_{O}\Delta_{A,b+1}$, $\Delta_{A,b+1}=e_{\text{eviction}}$ from (9).
(Case 2) Multiple attention heads: this is trivial and can be achieved by replacing $\Delta_{A}=\text{concat}(\Delta_{A}^{1},\dots,\Delta_{A}^{h})$, where super-script denotes head number.
(Case 3) Single self-attention and feedforward-network (FFN): we still assume single head without layer-norms. The dense logit is given as
$X^{1/2}_{b+1}=$ $X^{0}_{b+1}+W_{O}X^{0}_{A,b+1}$ (55)
$X^{1}_{b+1}=$ $X^{1/2}_{b+1}+W_{FFN}X^{1/2}_{b+1}$ (56)
$=$ $X^{0}_{b+1}+W_{O}X^{0}_{A,b+1}$
$+W_{FFN}X^{0}_{b+1}+W_{FFN}W_{O}X^{0}_{A,b+1}$ (57)
$l_{b+1}=$ $W_{H}X_{b+1}^{1}$ (58)
where, for simplicity we assume feedforward network to subsumed within $W_{FFN}$.
The perturbed logit due to eviction is given as:
$\hat{X}^{1/2}_{b+1}=$ $X^{0}_{b+1}+W_{O}X_{A,b+1}^{0}+{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}W_{O}\Delta_{A}^{0}}$ (59)
$\hat{X}^{1}_{b+1}=$ $\hat{X}^{1/2}_{b+1}+W_{FFN}\hat{X}^{1/2}_{b+1}$ (60)
$=$ $X^{0}_{b+1}+W_{O}X_{A,b+1}^{0}+{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}W_{O}\Delta_{A}^{0}}$
$+W_{FFN}X^{0}_{b+1}+W_{FFN}W_{O}X_{A,b+1}^{0}$
$+{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}W_{FFN}W_{O}\Delta_{A}^{0}}$ (61)
$\hat{l}_{b+1}=$ $W_{H}\hat{X}_{b+1}^{1}$ (62)
$=$ $l_{b+1}+{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\Delta_{l,b+1}}$ (63)
where, the logit error $\Delta_{l,b+1}=W_{H}(W_{O}\Delta_{A}^{0}+W_{FFN}W_{O}\Delta_{A}^{0})$. Thus, the above analysis shows that error in attention output can cause discrepancy in logit space which can affect performance on downstream tasks.
Note that for multiple layers, each layer would have its own eviction error which will keep compounding; however, this computing the exact compounded error is non-trivial due to the presence of output layer-norms.
Appendix E Additional Results
Layer 1 2 3 4 5 6 7 8 9 10 11 12 13 14
0.99 0.98 0.91 0.97 0.92 0.88 0.85 0.82 0.81 0.90 0.84 0.92 0.83 0.86
layer 15 16 17 18 19 20 21 22 23 24 25 26 27 28
0.91 0.88 0.94 0.96 0.92 0.95 0.99 0.99 0.96 0.98 0.98 0.88 0.99 0.96
Table 8: Spearman’s correlation between CAOTE and FastCAOTE on Llama3.2-3B-Instruct using sample from NarrativeQA
E.1 LongBench
Table 9: LongBench results for Llama 3.1-8B and Llama 3.2-3B-Instruct. Higher number is better. We highlight the best performing methods within a given budget with bold and the second best with underline.
Single Doc. QA Multi Doc. QA Summarization Fewshot Learning Synthetic Code
Narrative QA Qasper MF-en HotpotQA 2WikiMQA Musique GovReport QMSum MultiNews TREC TriviaQA SAMSum PCount PR-en Lcc RB-P Avg.
Llama 3.1-8B 30.05 47.00 56.12 57.33 47.81 32.25 34.86 25.32 27.02 73.00 91.61 43.37 8.33 99.50 61.66 51.94 49.20
6k Sink 25.41 47.40 44.13 47.39 45.73 21.90 32.53 22.19 26.87 72.00 91.25 43.41 3.08 52.50 62.22 56.24 43.39
H2O 8.52 43.31 44.80 40.03 42.46 21.68 11.85 8.78 26.03 62.00 56.39 25.72 5.75 45.50 58.62 29.53 33.19
+ CAOTE 25.77 46.45 54.99 50.57 47.7 31.93 33.99 22.86 27.01 71 87.05 41.29 6.67 54 60.48 43.23 44.06
+ FastCAOTE 27.02 46.46 55.4 51.32 47.4 32.89 34.08 23.69 27.03 71 86.25 42.14 9 48.5 60.49 42.94 44.10
TOVA 24.59 45.93 53.92 55.09 47.43 25.07 32.33 24.10 27.00 68.50 90.81 43.89 4.25 67.00 61.50 52.39 45.24
+ CAOTE 24.23 45.88 53.5 52.96 49.59 27.02 32.62 23.86 27.08 70 90.98 43.45 3 74.5 61.46 51.76 45.74
+ FastCAOTE 24.17 46.07 53.8 53.53 48.11 26.49 32.64 23.88 27.01 70 90.81 43.53 3 73 61.49 51.43 45.56
SnapKV 24.10 45.57 50.44 53.12 48.41 24.27 33.43 23.53 27.03 71.50 92.28 43.58 5.25 98.00 61.32 52.16 47.12
+ CAOTE 25.97 46.09 51.54 55.19 47.41 26.48 33.32 24 27.05 71.5 91.11 43.55 6.83 99.5 61.11 51.45 47.63
+ FastCAOTE 24.77 46.06 52.18 56.72 47.01 26.24 33.41 23.8 26.99 73 91.31 43.6 5.92 99.5 61.5 51.37 47.71
8k Sink 23.53 46.63 48.68 49.61 47.16 21.14 33.10 23.20 26.92 72.00 91.29 43.79 3.25 66.00 62.18 56.43 44.68
H2O 13.85 44.94 47.81 43.64 44.90 23.65 18.78 11.35 26.49 69.50 69.05 33.41 5.25 62.50 59.74 36.26 38.20
+ CAOTE 27.74 46.67 54.97 52.71 48.28 33.66 34.51 24.73 26.99 73 86.8 42.86 5 66.5 61.06 48.29 45.86
+ FastCAOTE 28.8 47.08 54.67 52.95 47.13 34.36 34.21 24.53 27.04 72 87.87 43.03 5.5 69.5 61.06 49.69 46.21
TOVA 24.86 46.78 54.83 54.52 49.00 26.40 33.44 24.76 27.00 71.00 91.11 43.29 6.25 87.00 61.49 51.79 47.09
+ CAOTE 25.65 46.88 54.5 55.42 48.73 26.54 33.47 24.8 27.02 72 91.11 43.26 6.25 87 61.36 51.3 47.21
+ FastCAOTE 25.25 46.75 54.76 56.29 48.94 26.25 33.37 24.81 27.01 71.5 91.11 43.24 5.25 89 61.32 52.1 48.58
SnapKV 25.15 46.55 53.39 56.00 48.75 27.82 33.67 24.85 27.01 72.50 91.78 43.54 5.08 100.00 61.48 51.41 48.06
+ CAOTE 27.06 46.42 53.76 56.51 47.79 28.13 33.87 24.93 27.02 73 91.38 43.29 6.75 99.5 61.49 51.99 48.31
+ FastCAOTE 26.91 46.59 53.47 56.63 48.54 29.27 33.91 24.86 27.01 73 91.38 43.49 6.75 100 61.49 51.78 48.44
Llama 3.2-3B 23.76 40.23 50.09 50.69 42.29 26.84 33.09 24.30 25.21 72.50 90.11 42.58 3.00 96.50 56.22 56.52 45.87
6k Sink 19.33 40.29 37.95 46.48 40.29 15.31 30.43 21.35 25.14 71.50 88.93 42.04 3.50 47.00 56.55 54.11 40.01
H2O 4.62 38.81 39.06 34.66 35.52 15.21 10.51 10.01 24.25 61.50 53.23 27.37 0.50 13.00 54.55 32.29 28.44
+CAOTE 16.14 41.68 49.36 46.7 43.36 22.75 32.07 21 25.07 69 80.02 39.33 1.5 26 55.82 49.05 38.68
+FastCAOTE 16.31 41.94 49.17 45.64 41.83 21.68 32.07 20.73 25.02 68.5 80.34 39.88 3.5 24 55.83 48.7 38.45
TOVA 20.22 39.78 45.86 49.08 41.54 20.43 30.50 22.17 25.11 66.50 89.00 42.50 4.00 46.50 55.57 57.53 41.02
+CAOTE 21.17 39.69 47.21 48.82 41.7 20.59 30.72 22.36 25.1 68 89 42.38 3.5 52.5 55.6 57.09 41.59
+FastCAOTE 21.48 39.66 47.02 47.56 41.95 19.91 30.8 21.98 25.17 67.5 89.5 42.06 4 53 55.6 57.39 41.54
SnapKV 20.83 39.65 44.48 49.30 40.18 20.28 31.27 22.73 25.09 69.00 89.95 41.47 4.00 85.00 55.69 57.82 43.55
+CAOTE 20.23 39.65 44.91 50.16 40.58 21.32 31.23 22.51 25.13 69 90 41.83 5 89.5 55.84 57.24 44.01
+FastCAOTE 20.09 40.02 44.58 48.57 42.12 22.51 31.25 22.89 25.15 71 90 41.83 4 89.5 55.83 57.25 44.16
8k Sink 20.15 40.02 41.94 48.15 42.24 16.01 31.64 22.10 25.20 73.00 89.26 42.37 3.50 62.50 56.86 56.63 41.97
H2O 9.65 39.66 43.20 38.09 40.41 21.46 17.80 13.28 24.67 70.00 64.30 32.19 2.00 24.50 55.00 39.09 33.46
+CAOTE 20.07 40.73 47.76 47.25 42.88 23.19 32.41 22.01 25.15 71 83.58 40.8 3 43.5 55.45 53.35 40.76
+FastCAOTE 20.81 40.54 48.1 47.35 43.4 25.13 32.73 22.31 25.18 71.5 84.91 40.6 4 45 55.84 52.89 41.27
TOVA 21.08 40.67 49.07 48.69 41.93 23.05 31.64 22.85 25.21 69.00 89.25 42.19 2.50 71.00 55.77 57.47 43.21
+CAOTE 21.97 40.66 49.37 50.1 41.29 24.05 31.65 22.85 25.16 69.5 89.5 42 3 78.5 55.82 57.16 43.91
+FastCAOTE 22.73 40.51 49.36 50.18 42.26 24.45 31.68 23.09 25.16 69.5 89.5 42.28 4.5 80 55.79 57.16 44.26
SnapKV 20.49 40.80 48.16 48.78 41.65 24.79 31.81 23.46 25.17 70.00 90.17 41.99 5.00 94.00 55.77 57.29 44.96
+CAOTE 19.71 40.7 48.05 49.03 41.27 22.95 31.95 23.1 25.21 72 90 41.88 4 95 55.77 57.02 44.85
+FastCAOTE 20.13 40.71 48.35 48.62 41.04 24.38 32.19 23.04 25.20 72 90 42.33 3.5 95 55.77 57.03 44.96
Table 10: LongBench results for Qwen 2.5-7B/2.5-3B-Instruct. Higher number is better. We highlight the best performing methods within a given budget with bold and the second best with underline.
Single Doc. QA Multi Doc. QA Summarization Fewshot Learning Synthetic Code
Narrative QA Qasper MF-en HotpotQA 2WikiMQA Musique GovReport QMSum MultiNews TREC TriviaQA SAMSum PCount PR-en Lcc RB-P Avg.
Qwen 2.5-7B 15.75 16.94 32.38 11.89 11.88 7.95 34.33 19.91 22.67 65.5 87.05 44.75 4.22 93.08 57.74 61.84 36.74
6k Sink 7.37 16.61 25.73 11.29 11.27 5.69 31.47 18.72 22.86 64.5 84.86 44.47 3.59 41.48 55.89 55.99 31.36
H2O 3.34 14.79 23.94 11.45 11.3 5.52 14.63 14.27 22.06 55.75 51.99 28.01 1.39 9.41 54.68 38.32 22.55
+ CAOTE 4.78 18.06 32.49 16.23 17.28 9.57 29.81 18.04 22.86 59.5 63.05 36.91 2.7 28.25 55.13 42.42 28.57
+ FastCAOTE 5.69 16.99 32.62 18.22 16.58 10.48 30.3 17.71 22.88 59.5 62.95 36.29 2.1 27.65 56.3 40.65 28.56
TOVA 15.77 15.33 30.31 19.3 13.78 9.11 30.4 19.95 22.91 61.5 83.47 42.9 1.15 21.775 57.68 57.99 31.46
+ CAOTE 15.81 16.07 29.39 19.4 14.15 10.8 30.89 20.54 22.86 62 84.92 43.19 2.17 30 57.76 57.53 32.34
+ FastCAOTE 15.67 16.23 30.4 19.45 13.32 10.18 30.77 20.2 22.82 61.5 83.29 43.3 1.5 28.75 57.71 58.1 32.07
SnapKV 14.34 16.35 31.12 17.56 14.1 8.74 31.09 20.16 22.84 60 83.8 42.99 2.91 54.17 57.48 60.26 33.62
+ CAOTE 12.77 16.3 31.33 19.74 14.06 11.07 31.02 20.85 22.91 61.5 83.79 42.97 4.8 68.25 57.54 61.08 35.00
+ FastCAOTE 12.98 15.93 31.3 18.58 13.82 9.45 30.96 20.27 22.88 61.5 84.58 43.28 5.34 65.48 57.54 60.25 34.63
8k H2O 6.1 15.55 28.29 12.37 14.65 6.24 20.78 17.22 22.44 59 58.74 33.05 1.82 15.73 55.63 44.56 25.76
+ CAOTE 8.65 15.59 34.92 20.41 15.95 13.6 32.11 20.05 22.82 63.5 78.2 40.66 3.85 46.33 57.19 51.7 32.85
+ FastCAOTE 6.76 15.88 34.3 20.75 16.2 16.82 31.95 20.67 22.81 62.5 77.33 41.02 3.03 47.08 57.23 50.48 32.8
TOVA 15.69 15.55 33.09 18.37 13.99 11.26 31.33 20.17 22.82 62 84.49 43.01 2.78 30.33 57.45 58.96 32.58
+ CAOTE 16.38 15.46 32.16 17.86 14.24 12.76 31.34 20.2 22.8 61 83.97 43.23 2.01 38.83 57.45 59.41 33.07
+ FastCAOTE 17.06 15.55 32.32 17.57 14.14 13.03 31.43 20.3 22.78 62 84.86 43.3 2.41 41.58 57.41 58.93 33.42
SnapKV 15.6 15.81 33.47 18.02 14.49 10.53 31.99 20.09 22.84 61 84.08 43.01 4.58 64.25 57.46 60.59 34.86
+ CAOTE 15.55 15.57 33.89 21.08 14.43 12.38 31.41 20.73 22.83 61.5 85.11 43.39 5.22 75.75 57.44 60.35 36.04
+ FastCAOTE 13.38 15.77 33.97 19.78 15.08 13.01 31.44 20.69 22.77 62 85.66 43.69 4.24 75.33 57.44 60.04 35.89
Qwen 2.5-3B 18.08 22.49 39.72 27.86 20.45 18.93 32.8 23.74 24.89 67.5 85.05 43.88 5 40.97 51.91 47.53 35.68
6k Sink 13.01 20.03 32.59 18.62 15.77 9.37 30.98 20.7 24.97 66.5 75.39 42.77 4 14.92 52.32 50.35 30.77
H2O 5.52 18.62 27.93 12.61 15.07 4.26 14.92 13.89 24.21 58 45.94 24.93 2.91 9.1 49.5 34.54 22.62
+ CAOTE 8.23 21.34 36.28 22.43 17.92 13.53 31.57 21.2 24.91 65 74.3 39.09 4.58 21.25 50.47 42.71 30.93
+ FastCAOTE 9.29 20.47 35.8 21.67 18.14 13.65 31.34 20.52 24.82 64.5 75.88 39.16 5.72 20.42 50.6 44.38 31.02
TOVA 13.62 19.56 34.64 21.67 16.25 8.47 30.17 23.1 24.94 63.5 81.88 42.97 1.16 10.58 51.3 47.7 30.72
+ CAOTE 13.28 19.82 35.25 22.6 15.5 8.97 30.4 23.17 24.84 64.5 81.68 43.46 2.07 13.21 50.56 47.05 31.02
+ FastCAOTE 13.34 19.71 35.58 22.02 15.65 8.76 30.49 23.26 24.82 65 80.92 43.66 1.75 13 51.45 47.74 31.07
SnapKV 14.16 20.09 36.15 19.14 15.59 12.7 30.35 22.75 24.91 65 83.92 43.52 5.00 32.2 51.04 47.49 32.75
+CAOTE 14.3 20.13 34.89 20.32 15.06 12.85 30.61 23.16 24.9 66.5 84.75 43.3 4.75 33.9 51.48 48.31 33.08
+FastCAOTE 14.33 19.48 34.59 20.8 16.54 11.85 30.68 23.19 24.93 66.5 83.54 43.55 4.62 33 51.06 47.93 32.91
8k H2O 6.16 19.84 32.32 16.01 17.74 4.99 20.21 16.49 24.54 64 56.1 32.56 3.13 11.61 50.61 38.8 25.94
+CAOTE 11.53 21.59 38.02 25.62 20.19 15.11 32.18 21.82 24.81 67.5 79.32 41.2 5.15 25.5 50.72 45.27 32.85
+FastCAOTE 11.65 21.2 37.92 24.47 20.32 13.25 32.11 21.75 24.84 67.5 78.07 40.27 5.17 25.21 50.17 46.37 32.52
TOVA 14.66 20.93 37.77 22.57 17.08 9.63 31.12 23.17 24.83 67 84.11 43.55 2.06 13.08 51.32 47.64 31.91
+CAOTE 13.98 21 36.91 22.97 17.05 9.93 31.25 23.45 24.9 66.5 84.14 43.86 3.07 13.92 51.14 47.883 32.00
+FastCAOTE 14.84 20.66 37.45 23.05 17.07 10.16 31.2 23.47 24.85 66.5 84.01 43.41 3.31 13.25 51.19 48.11 32.03
SnapKV 12.76 20.88 37.1 22.49 18.19 13.83 31.33 23.37 24.8 65.5 84.88 44.49 5.2 35.83 51.31 47.82 33.74
+CAOTE 13.66 20.41 38.08 24.76 17.31 13.21 31.3 23.62 24.82 66.5 84.88 44.16 5.17 36.58 51.24 47.94 33.98
+FastCAOTE 14.56 20.99 37.61 25.56 18.02 13.89 31.37 23.27 24.83 66.5 84.88 44.01 5 35.92 51.13 48.14 34.11
LongBench result for 66k and 88K for Llama 3.2-3B-Instruct/3.1-8B-Instruct and Qwen 2.5-3B-Instruct/8B-Instruct are shown in Table Table 9, Table 10 respectively. We also include Sink attention results (Xiao et al., 2024) with budget of 66k and 88k. For Llama 3.1-8B-Instruct, TOVA-FastCAOTE performs best for 66k budget, while SnapKV-FastCAOTE for 88k budget. For Llama 3.2-3B-Instruct SnapKV-FastCAOTE performs best for both 66k and 88k budget. On the other hand, for Qwen 2.5-7B-Instruct, SnapKV-CAOTE performs the best for both 66k and 88k, and for Qwen 2.5-3B-Instruct, SnapKV-COATE performs best for 66k budget, and SnapKV-FastCAOTE performs best for 88k budget. Additionally, note that all baseline token eviction methods achieve boost in accuracy when using CAOTE or FastCAOTE.
E.2 Perplexity
Table 11: Perplexity difference between different eviction methods with dense baseline. Lower is better. Negative entry in table means the method performs better than dense baseline. The PPL of Qwen 2.5-3B-Instruct and Qwen 2.5-7B-Instruct is 8.4547 and 7.3188 respectively.
Budget H2O TOVA SnapKV
+CAOTE +FastCAOTE +CAOTE +FastCAOTE +CAOTE +FastCAOTE
Qwen 2.5-7B-Instruct
2k 0.4253 0.4422 0.3917 0.0567 0.059 0.1077 0.0369 0.0987 0.0307
Qwen 2.5-3B-Instruct
2k 0.2585 0.2168 0.2154 0.0603 0.0513 0.0507 0.0278 0.0199 0.0196
We show perplexity results for Qwen 2.5 models in Table 11 for budget of 22k. SnapKV-FastCAOTE performs best for both Qwen 2.5-3B-Instruct and 2.5-7B-Instruct, and using CAOTE, all methods achieve improved perplexity.
E.3 Needle in a Haystack
Table 12: Needle-in-haystack accuracy for Qwen 2.5-3B/2.5-7B-Instruct using baseline eviction methods with(out) CAOTE. Higher is better, maximum accuracy is $1.0$.
Budget H2O TOVA SnapKV
+CAOTE +FastCAOTE +CAOTE +FastCAOTE +CAOTE +FastCAOTE
Qwen 2.5-7B-Instruct
6k 0.206 0.312 0.3 0.292 0.292 0.286 0.32 0.33 0.332
Qwen 2.5-3B-Instruct
6k 0.212 0.288 0.27 0.282 0.286 0.288 0.304 0.324 0.336
Table 12 shows Needle-in-haystack results for Qwen2.5 models for budget=66k. SnapKV-FastCAOTE performs best for both Qwen 2.5-3B-Instruct and 2.5-7B-Instruct, and using CAOTE, all methods achieve improved accurracy.
E.4 Ablations
Difference in Attention Outputs
Figure 3: Normalized error between attention outputs with and without eviction for Llama 3.2-3B-Instruct, measured on 5 samples from Qasper dataset.
In Figure 3, we check the average change in attention score in different layers for various eviction methods to provide insight into the performance gain realized by CAOTE. We measure the difference between the dense attention-output and eviction method based attention output, both with and without CAOTE. We repeat this for each layer and average the result to verify that CAOTE results in smaller $L_{2}$ distance to dense attention-output. Figure 3 shows the normalized mean squared error between the dense attention output and eviction-based attention output for H2O, TOVA and SnapKV, both with and without CAOTE in dashed-lines and solid-lines respectively. We observe that the CAOTE variant always produces a smaller gap from dense attention output (no eviction) compared to baselines, supporting our claim that CAOTE evicts the token which minimizes deviation from true attention output. Furthermore, we show that discrepancy in attention output due to eviction (eviction error) causes discrepancy in logits in D.2; to emphasize that optimizing for eviction error results in less discrepancy in logits and, therefore, less deviation on downstream task accuracy.
Different Prompt Block Sizes
Table 13: LongBench accuracy for different tasks from each sub-category with varying block-prompt-size using Llama 3.2-3B-Instruct model with 44k budget. $\infty$ implies entire prompt is consumed without token eviction and eviction starts during generation. We faced out-of-memory issues with some inputs. For $\infty$ case, HotPotQA gave out-of-memory on a 80GB A100 GPU and ran only on 114 out of 200 samples.
Task H2O H2O-CAOTE TOVA TOVA-CAOTE SnapKV SnapKV-CAOTE
block-size 32 128 $\infty$ 32 128 $\infty$ 32 128 $\infty$ 32 128 $\infty$ 32 128 $\infty$ 32 128 $\infty$
HotPotQA 20.43 24.49 35.71 32.37 40.28 46.37 44.36 45.55 52.93 44.72 45.25 52.94 45.99 46.7 53.67 44.66 45.64 52.67
Qasper 30.8 31.94 32.58 36.74 40.79 38.93 38.65 38.83 40.45 39.39 39.79 39.90 38.71 39.22 39.80 40.18 39.74 39.69
GovReport 5.52 5.44 7.07 30.10 30.82 31.48 29.32 29.41 31.25 29.24 29.39 31.15 29.62 29.79 31.82 30.15 30.07 32.42
Trec 46.50 53.00 56.00 59.00 65.5 72.00 62.00 64.00 71.50 64.00 63.00 71.50 64.00 65.50 71.50 67.50 67.50 71.50
LCC 51.59 51.23 51.51 54.06 55.32 54.51 56.23 55.64 56.34 56.18 55.68 56.39 56.24 55.74 56.56 56.27 55.73 56.59
[…]